Justify - H*(CO2) = Delta_f H*(CO2) [*means zero in superscript and _f means f in subscript]

where, H* is standard enthalpy and Delta_f H* means standard enthalpy change of formation.
1
Expert's answer
2014-06-10T11:04:13-0400
C + O2 --> CO2

Delta_f H*(CO2) = H*(CO2) - [H*(C) + H*(O2)]
The standard enthalpy of elementary substance is zero:
H*(C) = H*(O2) = 0
Therefore
Delta_f H*(CO2) = H*(CO2) - [H*(C) + H*(O2)] = H*(CO2) - [0 + 0] = H*(CO2)